Dublin University Rifle Club

Version 1.2 Date : May 18, 2001 Filename : /home/dennehym/DURC/Reports/RCS/CorrespondanceRecord.tex,v Status : Exp

Secretary/Match Secretarys Correspondance Record

May 21, 2001

1 Mail Received

1 Mail Received

Date From  Description
15 May 2001 Bob Green  UCESSA 50m/100yds championships details

2 Mail Sent

2 Mail Sent

Date To  Description

16 May 2001 Competition Attendance Records list Competition Attendance Reports on Air Nationals and Prone Nationals, accompanied by a cover letter explaining same.

16,17 May 2001 RO Dinner List  Letter asking for suggested dates for RO Dinner

17 May 2001 AGM Distribution list Copy of AGM minutes of May 9th AGM. Attached : Captains Report, Treasurers Report, Chief Range Officers Report, Match Secretary Report, 1993 Constitution, proposed changes to same, 2001 Constitution (with changebars)

18 May 2001 P. Gray  NSRA/Eley Postal cards, Round 2.

via Registered Mail 8 cards by Ben Jones. Est Scores 42,30,42,43,39,39,41,41. Witnessed by Mark Dennehy, dated 14 May. Scanned images stored in dunsinane. Also enclosed, a note regarding the receipt of only four stickers for the eight cards.

18 May 2001 B. Wingrove  NSRA/Eley Postal cards, Round 2.

via Registered Mail 4 cards by Ray Reilly. Witnessed by John Keeney, dated 14 May. Est Scores [96-94], [99-98], 100(9X), [97-96] 4 cards by John Keeney. Witnessed by Ray Reilly, dated 13 May. Est Scores [99-98], 94, [98-96], [96-94] 4 cards by Alan Smyth. Witnessed by Mark Dennehy, dated 14 May. Est Scores [100-97], 95, [95-93], [96-92] Scanned images stored in dunsinane.

18 May 2001 Geoff Cooney Note about the TTT and how I’d like to see him get 15-20 juniors to sign up at a pound a head

email

18 May 2001 Gerry Power  Note about TTT and would he get some austrailians to enter

email

21 May 2001 TTT distribution list Notice letters regarding upcoming TTT. Personal letters to all clubs on list outside Ireland.

21 May 2001 Adrianne Regan, Head of Sponsorship, Eircom

Initial Sponsorship Proposal letter

Registered Post

21 May 2001 College post All mail sent from 18 May on returned. College will not post our correspondance. Stamps obtained and all mail posted before 1800 today.

3 Mailing List Details

3 Mailing List Details

Competition Attendance Records distribution list

DURC President Terry McAuley
Director of Sport Terry McAuley
DUCAC Chairperson Dr. Trevor West
DUCAC Administrative OfficersCathy Doyle

Geraldine McAuley

Physical Recreation Officer Michelle Bennett
Sports Facilities Committee Helen Murphy

AGM Distribution list

DURC President Terry McAuley
Director of Sport Terry McAuley
DUCAC Chairperson Dr. Trevor West
DUCAC Administrative OfficersCathy Doyle

Geraldine McAuley

Physical Recreation Officer Michelle Bennett
Sports Facilities Committee Helen Murphy

RO Dinner list

Mark Dennehy  John Egan  Dave Cooney
Geoff Cooney  Rhona Barry  Gary Duff
Ray and Collette Kane Alan Lewis  Pat O’Brien
Gerry Power  Nicola Murphy  Paul O’Boyle
Barry Dagger  Albert Jordan  Ian Marnane
Aisling Broderick  Alan Smyth  John Keeney
Ray Reilly  Jason Sibley  Ben Jones
Sharon Wall  Michael Madigan Shane Kelly
David Gallagher  Simon Kennedy